# The AI Productivity Index: APEX-v1-extended

### v3

@@ 1 Introduction
Changes to grading in APEX-v1-extended include: (1) switching the panel of LM judges to a single LM, Gemini 2.5 Pro (Thinking = On), (2) passing the prompts to each model 8 times and taking a mean (rather than 3 times and taking a median) and (3) reporting 95% confidence intervals with model scores. We have used this methodology to create a new leaderboard for APEX-v1-extended with results for 10 frontier models.333mercor.com/apex. GPT 5 (Thinking = High) has the highest mean score at $67.0\%$, followed by Gemini 3 Pro (Thinking = High) at $64.3\%$ and Grok 4 at $63.5\%$. APEX-v1-extended shows that models struggle with real-world tasks that professionals undertake everyday.
@@ 2.2 Quality control
Experts created each prompt independently, based on their day-to-day work. The mean time for an expert to execute the prompts in a real-world professional setting is estimated at $2.7$ hours, with a range of $0.5$ to $20$ hours. Experts submitted their prompts for review, where they were approved or rejected by reviewers. For rejections, reviewers could either remove the prompt from production or request specific changes. After the prompt was approved, the contributor created the grading rubric, which was also reviewed by a reviewer. Multiple rounds of review help to ensure that cases meet the requirements of the project, are varied in terms of topic, style and model behavior, and are sufficiently challenging. We use in-house LM-powered reviewing tools to give experts immediate feedback on their work, although experts are always responsible for the quality of their submissions. The quality control process did not change from APEX-v1.0 to APEX-v1-extended.
@@ 3.1 Model selection
We tested 10 frontier models from Anthropic, Google DeepMind, OpenAI, and xAI on APEX-v1-extended. Responses were collected at the end of November 2025 from the models’ respective APIs. Thinking is turned On if available and set to High if configurable. If temperature can be configured, we set it to $0.7$. We do not explicitly set the system prompt. For each prompt, we collect model responses 8 times. This is an increase from 3 times in APEX-v1.0, allowing us to calculate the mean (rather than median) and confidence intervals.
@@ 3.2 Grading responses with an LM judge
LM judges are widely used to scalably assess the quality of model responses, supporting high-fidelity evals that can be re-run with no marginal human work (Baumann et al., 2025; Gu et al., 2025; Zhu et al., 2025). LM judging is well-suited for grading rubrics as each criterion is a short self-contained statement. In principle, these statements are easier to judge than overall quality or abstract concepts like “usefulness” or “helpfulness”. We used a single judge in APEX-v1-extended (Gemini 2.5 Pro, Thinking = On, Temperature = $0.1$) rather than the panel of judges for APEX-v1.0. This reduces the eval time, is more transparent, and maintains high quality. The prompt used by the LM judge remains the same in APEX-v1-extended as APEX-v1.0, and is given in Appendix A.
@@ 4.3 Comparison with APEX-v1-devset
We evaluate the same 10 models against the $n=100$ cases in APEX-v1-devset (available open source) using the exact same grading methodology (i.e., 8 runs and Gemini 2.5 Pro (Thinking = On) as a judge). Results are shown in Table 4. Overall, the open source data is similar but slightly more challenging. GPT 5 (Thinking = High) is the best scoring model on both sets, achieving $65.70\%$ on the APEX-v1-devset versus $67.0\%$ on APEX-v1-extended (difference = $1.3\%$). Differences in model ranks are minor (two sets of two models are swapped one position) and only one model performs better on the dev set than the benchmark (Gemini 3 Pro (Thinking = High) scoring $1.2\%$ higher). Most models find APEX-v1-devset harder, with GPT 5.1 (Thinking = High), Gemini 2.5 Flash (Thinking = On) and Sonnet 4.5 (Thinking = On) scoring lower by at least 10 percentage points.Finally, over the 8 runs, the mean standard deviation for APEX-v1-devset is slightly higher, increasing from $9.05$ to $10.76$.
@@ Appendix A LM judge
Implementation details: We use Gemini 2.5 Pro as a judge LM, with temperature set to 0.1 and Thinking turned on.
A.1 LM judge prompt

### v4

@@ 1 Introduction
Changes to grading in APEX-v1-extended include: (1) switching the panel of LM judges to a single LM, Gemini 2.5 Flash, (2) passing the prompts to each model 8 times and taking a mean (rather than 3 times and taking a median) and (3) reporting 95% confidence intervals with model scores. We have used this methodology to create a new leaderboard for APEX-v1-extended with results for 10 frontier models.333mercor.com/apex. GPT 5 (Thinking = High) has the highest mean score at $67.0\%$, followed by Gemini 3 Pro (Thinking = High) at $64.3\%$ and Grok 4 at $63.5\%$. APEX-v1-extended shows that models struggle with real-world tasks that professionals undertake everyday.
@@ 2.2 Quality control
Experts created each prompt independently, based on their day-to-day work. The mean time for an expert to execute the prompts in a real-world professional setting is estimated at $2.7$ hours, with a range of $0.5$ to $20$ hours. Experts submitted their prompts for review, where they were approved or rejected by reviewers. For rejections, reviewers could either remove the prompt from production or request specific changes. After the prompt was approved, the contributor created the grading rubric, which was also reviewed by a reviewer. Multiple rounds of review help to ensure that cases meet the requirements of the project, are varied in terms of topic, style and model behavior, and are sufficiently challenging. We use in-house LM-powered reviewing tools to give experts immediate feedback on their work. Experts are always responsible for the quality of their submissions. Finally, to ensure that models give a valid single-turn response, we append a short instruction to each prompt, as given in Appendix B.
@@ 3.1 Model selection
We tested 10 frontier models from Anthropic, Google DeepMind, OpenAI, and xAI on APEX-v1-extended. Responses were collected at the end of November 2025 from the models’ respective APIs. Thinking is turned On if available and set to High if configurable. If temperature can be configured, we set it to $0.7$. For each prompt, we collect model responses 8 times. This is an increase from 3 times in APEX-v1.0, allowing us to calculate the mean (rather than median) and confidence intervals.
@@ 3.2 Grading responses with an LM judge
LM judges are widely used to scalably assess the quality of model responses, supporting high-fidelity evals that can be re-run with no marginal human work (Baumann et al., 2025; Gu et al., 2025; Zhu et al., 2025). LM judging is well-suited for grading rubrics as each criterion is a short self-contained statement. In principle, these statements are easier to judge than overall quality or abstract concepts like “usefulness” or “helpfulness”. We used a single judge in APEX-v1-extended (Gemini 2.5 Flash) rather than the panel of judges for APEX-v1.0. This reduces the eval time, is more transparent, and maintains high quality. The prompt used by the LM judge remains the same, as given in Appendix A.
@@ 4.3 Comparison with APEX-v1-devset
We evaluate the same 10 models against the $n=100$ cases in APEX-v1-devset (available open source) using the exact same grading methodology (i.e., 8 runs and Gemini 2.5 Flash as a judge). Results are shown in Table 4. Overall, the open source data is similar but slightly more challenging. GPT 5 (Thinking = High) is the best scoring model on both sets, achieving $65.70\%$ on the APEX-v1-devset versus $67.0\%$ on APEX-v1-extended (difference = $1.3\%$). Differences in model ranks are minor (two sets of two models are swapped one position) and only one model performs better on the dev set than the benchmark (Gemini 3 Pro (Thinking = High) scoring $1.2\%$ higher). Most models find APEX-v1-devset harder, with GPT 5.1 (Thinking = High), Gemini 2.5 Flash (Thinking = On) and Sonnet 4.5 (Thinking = On) scoring lower by at least 10 percentage points. Finally, over the 8 runs, the mean standard deviation for APEX-v1-devset is slightly higher, increasing from $9.05$ to $10.76$.
@@ Appendix A LM judge
Implementation details:
@@
Appendix B Prompt adjustment
Many models are trained to ask followup questions or outline options that the user can confirm or reject. To ensure they give a complete response, we append the following text to every prompt.
**Instruction**
You are an AI assistant that produces final, domain-appropriate deliverables from a given task description and (optionally) attached files. You will be given the following inputs:
  • Task Domain: Domain
  • Task Prompt: Prompt
  • Attachments:
==== Attached files content: ====
  === <File_1> ===
  <File_1_Contents>
  === <File_2> ===
  <File_2_Contents>
  … (repeat as needed)
Ground Rules
  1. 1.
    You must not ask follow-up questions. Interpret the prompt as best you can and produce the best complete answer given the information provided.
  2. 2.
    Use the attachments as primary sources.
  3. 3.
    Treat each "=== <File_Name> === …" block as the full content of that file.
All of the source files that you need have been added to the prompt.